OpenStudy (anonymous):

Fill in the blank. A pizza with a diameter of 16 inches is cut into 8 equal slices. Each slice of the pizza has an area of ______ square inches. Use your calculator and round your final answer to two decimal places.

OpenStudy (anonymous):

HELP (GIVEN AWAY MEDALS)

OpenStudy (anonymous):

Circumference = 2 • π • radius = π • diameter Circle Area = π • r² = ¼ • π • d²

OpenStudy (anonymous):

Circle Area = ¼ • π • d² Circle Area = ¼ • π • 16² Circle Area = ¼ • π • 256 Circle Area = 64 • π Circle Area = 64π

OpenStudy (anonymous):

64 IS THE FINAL ANSWER CORRECT?

OpenStudy (anonymous):

\[\frac{ 64 \pi }{ 8 } = 8\pi\]

OpenStudy (anonymous):

\[8\pi = 8 \times 3.14 \approx 25.12\]

OpenStudy (anonymous):

25.12 IS THE FINAL ANSWER